A fixed mass of gas of volume 600 cm^3 at a temperature of 27 degree celcius is cooled at constant pressure to a temperature of 0 degree celcius, what is the change in volume?
1
Expert's answer
2020-07-03T10:00:52-0400

According to the Charles's law, when a fixed mass of gas is is cooled at constant pressure, its temperature "T" and volume "V" obey the following equation:


"\dfrac{V}{T} = k"

Let's find the constant "k". In the initial state:


"k = \dfrac{V_1}{T_1} = \dfrac{600\cdot 10^{-6}m^3}{(273+27)K} = 2\cdot 10^{-6} m^3/K"

After cooling, the new volume will be:


"V_2 = kT_2 = 2\cdot 10^{-6} m^3/K\cdot 273K = 546\cdot10^{-6} m^3 = 546 cm^3"

Answer. 546 cm^3.